A researcher studying the nutritional value of a new candy places a 6.10-gram sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increase is 2.80 °C. If the heat capacity of the calorimeter is 30.60 kJ·K–1, how many nutritional Calories are there per gram of the candy?
1
Expert's answer
2015-10-08T04:18:50-0400
1 Cal = 1 kcal = 4.1868 kJ
Q = Ccalorimeter × ΔT;
Ccandy = Q/m = Ccalorimeter × ΔT/m = 30.60 kJ K-1 × 2.80 K/6.10 g = 14.0 kJ/g = 3.35 kcal/g = 3.35 Cal/g

Answer 3.35 Cal/g